OpenStudy (anonymous):

whats the domain of √x+√(6-x)?

OpenStudy (anonymous):

the number inside the radical must be greater than or equal to 0.. So you do this \[6 - x \ge 0\] solve for x \[x \le 6\] So the domain is \[(-\infty,6]\]